10 / (3-i2)
(10 divided by 3-i2)

Answer by mukhopadhyay(490)

10/(3-i2)
Our goal is to eliminate any radical or non-real complex numbers from the denominator. We can do this by multiplying both numerator and denominator of the given expression with (3+i2).
This gives [10/(3-i2)](3+i2/3+i2)
= 10(3+2i)/[3^2-(i2)^2]
= 10(3+2i)/(9+4)
= 10(3+2i)/13